For the discrete distribution given in the table below
find the probability that x is greater than 1: P(x>1)
x 0 1 2 3 4
P(x) 0.1 0.2 0.3 0.3 0.1

Answer by ewatrrr(24785)


Hi
x 0 1 2 3 4
P(x) 0.1 0.2 0.3 0.3 0.1
P(x > 1) = 1 - P(x<= 1) = 1 -(.1+.2) = .7